VIM spectrograms showed and increased rate of broadband magnetic glitches inside the CEB hall, which occurred since last week: Figure 1. Also seen by the IPS VOLT probe: Figure 2.
The signature reminded us of the noise from the DET HVAC humidifier encoutered 4 years ago (!): https://logbook.virgo-gw.eu/virgo/?r=53672.
As reported in Figure 1, the glitches disappeared when one of the two humidifiers (electrolitic type) got stuck on Sun. Oct 19. On Tuesday morning this humidifier helectrodes were cleaned (with time electrodes get covered by limestone and surge larger currents). The glitches rate reduced but did not disappear.
In recent times Roberto Cavalieri implemented current monitors of both humidifiers (named INF_CEB_HUM_1 (main) and INF_CEB_HUM_2 (spare)). These sensors help to charaterize the humidifiers noise.
Thanks to the monitors, we can tell that some glitches from the humidifiers (at least one every few minutes) are always present, and seen both in magnetometers and Voltage / Current IPS monitors of CEB (Figure 3). Also, according to the monitors, since Tuesday intervention both DET humidifiers are active (Figures 4 - September, Figure 5 - October, Figure 6 - October full).